OpenStudy (anonymous):

simplify -10+2i/2, -10-2i/2

OpenStudy (anonymous):

so two separate terms and no plus or minus sign between them?

OpenStudy (anonymous):

I meant to post the whole problem, sorry.t^2+10t+26=0 = -10± -10^2 -4*1*26/2*1 = -10 100-104/2 = 10± -4/2 t= -10+2i/2, -10-2i/2

OpenStudy (anonymous):

i need to finish simplifying. I thought it was but apparently not.